Ace Company manufactures two products, X and Y. The following information is given:

     Data regarding the two products follow:

Product

Direct labor hours required

to make 1 unit

Annual production units

X

4 hours per unit

2,000 units

Y

5 hours per unit

1,000 units

                                                                                                                                                           

     Data regarding direct material and direct labor costs follow:

                                      X                                  Y

Direct material cost per unit of product              $90                              $200

Direct labor cost per unit of product                     80                              100

                                                       

    This company uses activity-based costing. The company’s overhead costs are traced to the following two

   activity cost pools. The following data are given:

                                                            

Activity

cost pool

Activity measure

Estimated Overhead Cost

Activity

For X

For Y

Total

Part

Management

Number of parts

$12,000

200 parts

600 parts

800 parts

Special Processing

Machine hours (MHs)

$18,000

500 MHs

400 MHs

900 MHs

    Activity rates are already computed as follows:

                        Part management: $15/part

                        Special processing: $20/MH

10. How much overhead cost is assigned to Product X (for all 2,000 units together)? $_________

11. How much is the overhead cost per unit of Product X? $_________

12. How much is the unit product cost of Product X? $_________
Answer #1

10.Overhead Cost assigned = Activity Consumption*Activity Rate

= 15*200 + 20*500 = $13,000

11.Overhead cost per unit = Total Overhead Cost/Total Units

= 13,000/2,000

= $6.5

12.Unit product cost = Direct Material + Direct Labor + Overheads

= 90+80+6.5

= $176.5